# On the Robustness of Verbal Confidence of LLMs in Adversarial Attacks

### v1

@@ 6 Conclusion and Limitations
In conclusion, we have established the threat to LLMs from VCAs through a comprehensive study across different datasets, CEMs, and threat vectors. By formulating four perturbation-based and two jailbreak-based methods, we reveal serious vulnerabilities that can cause frequent and high levels of confidence score changes, along with answer changes as with conventional adversarial attacks. Through extensive analysis on the behaviour and properties of CEMs under a variety of attacks and mitigation strategies, we demonstrate that LLMs tend to maintain a high level of confidence irrespective of perturbations, and that current defences are not fully adequate at mitigating the effects of attacks. Ultimately, as black-box LLMs become more widely adopted among both general users and in safety critical tasks where getting uncertainty estimates is key, we believe a higher focus needs to be ascribed not only to creating new and better calibrated black-box probability measures utilizing verbal confidence, but ensuring that such measures are also robust by developing strategies for mitigating the ease with which they can be utilized for adversarial attacks and how simply they can be manipulated through trivial changes in input.
The nature of running adversarial attacks requires many LLM queries, so we are unable to run more extensive experiments on larger language models. Nevertheless, we highlight two powerful LLMs at different scales, and corroborate our results using two SOTA models. In addition, we primarily focus on character and word-level VCAs and have shown their effectiveness comprehensively. Sentence-level attacks can be constructed using a similar framework which we leave for future work. Lastly, we examine VCAs on reducing generated confidence, which mirrors the objective of traditional adversarial attacks. Attacks for increasing confidence are possible, but will have less room for attack effectiveness due to the high performance and confidence of victim LLMs.
@@ Appendix A Motivation
Firstly, we stress that studying verbal confidence and its robustness is an important topic given that numerous companies in industry have developed models that utilize verbal confidence in their methods to conduct confidence estimation given limitations such as black-box API access, and are shown to ultimately be concerned about robustness [7, 39, 71] which we will highlight. Chen and Mueller [7] develop a Trustworthy Language Model (TLM) where Reliability and explainability is added to every LLM output and decision-making is conducted using trustworthiness scores. This score quantifies the certainty of the model to its answers by combining checking the consistency and the level of verbal numeric confidence of the model’s responses. Zhang et al. [71], in the context of cloud incident root cause analysis, develop a method for on-call engineers to make decisions on whether to adopt model predictions using a method that incorporates an LLM-based confidence estimator using verbal confidence due to the black-box nature of the APIs of SOTA models. In MacDonald [39] they build an LLM-based product to structure unstructured data and score customer conversations for developing sales and customer support teams. These scores are based on level of customer satisfaction and given relevant (proprietary) context, the LLM produces a score (verbally) and reasoning for that score. In all of these use-cases it is clear that compromising the robustness of these verbal scores creates a risk of system malfunction or a degradation in operation.
@@
Many prior works such as Dhuliawala et al. [9], Zhang et al. [72], Rechkemmer and Yin [53], have established how harmful miscalibrated (skewed) confidence scores are in human studies and downstream human-AI collaboration applications, showing that user behaviour changes greatly in response to AI confidence scores, leading to over reliance when an AI system is overconfident, as one example. Most notably, Dhuliawala et al. [9] shows in the context of human collaboration with LLMs how un-confidently correct stimuli reduce user trust with only a few miscalibrated instances creating long-term poor performance on tasks.
Moreover, works such as Bernsohn et al. [3], Beam and Kompa [2], Penso et al. [50], Jiang et al. [27] highlight the impacts of altered confidence scores on downstream authority mechanisms in legal and medical domains, since many such applications that involve tasks such as diagnosis rely on model uncertainty to determine further review. In Bernsohn et al. [3] as an example, in the task of detecting legal violations, entities with high confidence undergo additional processing, and therefore reducing the confidence estimates in the examples would compromise the operation of the system.
@@
In addition to harm done by reduced confidence scores (when answers remain unchanged), our key contribution is also to demonstrate verbal confidence can be used to effectively craft conventional adversarial examples (i.e., ones that lead to answer changes) as well, which presents issues in practical scenarios (e.g., misclassified cases). Our work to our knowledge is the first to use verbal confidence scores as the attack objective in contrast to previous works. This is important because such scores are easy to obtain for adversaries for black-box models, and can be used to manipulate the confidence measures that can most easily be asked for and understood by human users.
@@
Our formulation of this issue also has contributions in terms of its implications towards model honesty and how models express their level of knowledge/competence through confidence. A key aspect of this study is to determine how easy their behaviour is to manipulate through various types of (mostly superfluous) modifications, and through our analysis determining among which aspects there are the most vulnerabilities. In short, one of our objectives is to see how models “perceive” confidence overall and hence we look at multiple different aspects of how models react to different changes of input. We primarily test generalist models to get a broad view of the problem of robustness and to avoid making any conclusions that would only apply to a singular domain.
@@
We study 3 different threat vectors. An adversary can also target demonstrations or the system prompt too without a user being aware when they send their own unmodified query. This can be a form of data poisoning attack since, for example, adversarially modified demonstrations could be used by a developer of an LLM system unknowingly from a database of seemingly benign demonstrations.
@@ Appendix B Details of Perturbation-based Attacks
The general algorithms for the four different character and word-level adversarial attack methods we cover can be seen in Algorithm 2. For VCA-TF and VCA-TB we make modifications on the black-box algorithms to make the algorithm work in the context of attacking confidence scores generated using CEMs, as these algorithms are originally based on using the predicted softmax class-wise probabilities derived from model logits. Refer to more details in Jin et al. [28] and Li et al. [33] respectively for more detailed settings as we are faithful to the original algorithm except with respect to the scoring functions used. For the Typos and SubSwapRemove attacks, we make use of our own design choices which we detail below. In the algorithm, the potential options for the target attack text $\mathbf{X}$ are user queries, system prompts, or demonstrations. Additional input text $\mathbf{Q}$ refers to the two remaining text inputs that are not undergoing the attack (e.g., if the user query is targeted, $\mathbf{Q}$ represents the system prompt and demonstration). The parts that are attacked for user queries refer to the question only, without the multiple choice options. For system prompts the entire prompt detailing the task instructions (i.e., how to generate answer and confidence score) can be attacked. For the demonstration attacks, the demonstration is a one-shot example for a given CEM. Within the example, the question, multiple choice options, and the expected answer can be targeted.
For VCA-TB, the possible random bugs that can be generated include inserting spaces into words, deleting characters, swapping adjacent letters, swapping characters with visually similar characters, and substituting words with one of its nearest neighbours. We use WordNet [12] to search for synonyms for each token and calculate the semantic similar using Euclidean distance with GloVe embeddings [49]. Likewise, with VCA-TF, we use the same approach to generate a list of synonyms, although we also ensure that synonyms contain the same part of speech (POS) tags using NLTK [38].
To perform the Typos attack, we generate a series of common typos in a sentence. These typos include substitutions with nearby characters on an English keyboard, random character deletions, and adjacent character swaps. A probability of $0.1$ is used to select the percentage of characters in a full text sequence to modify each iteration. We cap the number of iterations at 25.
@@
1:Input: Confidence estimation method $z$ for LLM, Tokenized Target Text $\mathbf{X}$, Additional Input Text $\mathbf{Q}$ Sentence Similarity Model SIM(), Perturbation Similarity Rejection Threshold $\epsilon$, Max Iterations $\tau$, # Synonym Candidates $N_{syn}$
@@
3:$\mathbf{\hat{X}}$leftarrow $\mathbf{X}$
4:$\mathcal{C}$leftarrow $z(Q,\mathbf{X})$triangleright Original Confidence Score
@@
9:     $\mathbf{J}_{\hat{x}_{i}}$ $=$ $\mathcal{C}-z(\mathbf{Q},\mathbf{\hat{X}}_{\backslash\hat{x}_{i}})$triangleright $\mathbf{\hat{X}}_{\backslash\hat{x}_{i}}$ is set $\mathbf{\hat{X}}$ without token $\hat{x}_{i}$
@@
15:     $Bug$leftarrow Randomly Selected TextBugger Bug on $w_{i}$
16:     $\mathbf{\hat{X}}_{temp}$leftarrow $\mathbf{\hat{X}}_{w_{i}<>Bug}$triangleright Bug replaces original token in $\mathbf{\hat{X}}$
[…]
@@
19:              $\mathbf{\hat{X}}$leftarrow $\mathbf{\hat{X}}_{temp}$
@@
[…]
28:         $\mathbf{\hat{X}}_{temp}$leftarrow $\mathbf{\hat{X}}_{w_{i}<>cnd}$triangleright cnd replaces $w_{i}$ in $\mathbf{\hat{X}}$
[…]
@@
31:                  $\mathbf{\hat{X}}$leftarrow $\mathbf{\hat{X}}_{temp}$
32:                  topscore $=$ $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})$
@@
43:              $\mathbf{\hat{X}}$leftarrow $\mathbf{\hat{X}}_{temp}$
@@
52:     $\mathbf{\hat{X}}_{temp}$leftarrow MOD$(\hat{x}_{i},\mathbf{\hat{X}})$
[…]
@@
55:              $\mathbf{\hat{X}}$leftarrow $\mathbf{\hat{X}}_{temp}$
@@ Appendix C Detailed Experimental Settings
Dataset Details: (1) MedMCQA [47] is a multiple-choice question answering (MCQA) dataset consisting of 2816 questions with four choices that are based on medical entrance exam questions. (2) TruthfulQA [36] is a MCQA dataset consisting of questions that cover a broad range of safety-critical topics like health, law, finance and politics and is designed so that humans can easily answer incorrectly due to misconceptions. We use the MC1 (Single-true) subtask, and only use multiple choice questions with 4 total answers for consistency yielding 203 total examples. (3) StrategyQA [16] consists of 2,780 true or false open-domain, general topic questions that require implicit reasoning steps to answer. Lastly, we use SST-2 [57] from the popular AdvGLUE [61] adversarial benchmark for our baseline confidence attack agnostic results. In the benchmark, SST-2 consists of 137 movie reviews annotated for binary (positive/negative) sentiment. Instances in AdvGLUE underwent a comprehensive series of word-level, sentence-level, adversarial algorithmic modifications, in addition to human-generated perturbations to most effectively deceive a wide range of language models. For our experiments, we use the development sets for MedMCQA and TruthfulQA, and the training set for StrategyQA.
Model Details: Llama-3-8B [1] and GPT3.5-turbo [45] are primarily used for evaluation, with GPT-4o [45] and Llama-3-70B [1] being utilized for certain additional results. The Llama models are ran locally while the GPT models are ran through the OpenAI API. In terms of sampling parameters for each model, we utilize a temperature of 0.0 for Base, CoT, and MS method, along with a top_p of 0.95. For SC, we set the temperature to 0.7 and sample three different responses and average the confidence scores. We limit the number of generated tokens to 400. For the MS method, we calculate the geometric mean of the generated confidence scores for each step including the final.
Sentence Similarity: To ensure adversarial examples preserve the original meaning, we encode both the original and adversarial example using Universal Sentence Encoder (USE) [6]. We then calculate cosine similarity between the vectors and forgo perturbations that cause a fall below a threshold of 80% similarity to keep adversarial examples retaining similar meaning of the original sentences similar to the approach in Wang et al. [63].
[…]

### v3

@@ 6 Conclusion and Limitations
We have established the threat to LLMs from VCAs through a comprehensive study across different datasets, CEMs, and threat vectors. By formulating four perturbation-based and two jailbreak-based methods, we reveal serious vulnerabilities that can cause frequent and high levels of changes in verbal confidence and predicted answers. Through extensive analysis on the behaviour and properties of CEMs under a variety of scenarios and mitigation strategies, we demonstrate inadequacies with existing defences at fully mitigating the effects of attacks. Ultimately, as black-box LLMs become more widely adopted among both general users and in safety critical tasks where getting uncertainty estimates is key, we believe a higher focus needs to be ascribed not only to creating new and better calibrated black-box probability measures utilizing verbal confidence, but ensuring that such measures are also robust. This can be done by developing strategies for mitigating the ease with which they can be utilized for adversarial attacks and how simply they can be manipulated through trivial changes in input. In this sense, future methods need to be robust against VCAs but not so robust that they prevent LLMs in responding adequately to situations where the confidence should change.
The nature of running adversarial attacks requires many costly LLM queries, so we are unable to run more extensive experiments on larger language models. Nevertheless, we highlight two powerful LLMs at different scales, and corroborate our results using recent SOTA models. In addition, we primarily focus on character and word-level VCAs and have demonstrated their effectiveness. Sentence-level attacks can be constructed using a similar framework which we leave for future work. Our main VCA objective is also confidence reduction to mirror conventional adversarial attacks. VCAs aimed at increasing confidence are possible, but leave less room to study attack effectiveness due to the high confidence of victim LLMs. Lastly, we note that the datasets we test possess low ecological validity [devic2025calibrationcollaborationllmuncertainty]. Nevertheless, we have focused on well-studied datasets which are widely used in previous works. This is important, as due to the inherent variability of real-world data it is important to explore the idea of VCAs with well-tested datasets in tasks that previous work has covered for ease of interpretability. We strongly encourage future work to explore how the unique properties of different domains and tasks affect VCAs.
@@ Appendix A Motivation
Firstly, we stress that studying verbal confidence and its robustness is an important topic given that there are numerous instances in industry where systems have been developed that utilize some form of LLM verbal confidence in their methods to conduct confidence estimation. They utilize verbal confidence largely due to limitations with black-box API access and since the scores are not arbitrary [pmlr-v235-mohri24a, Omar2024.08.11.24311810] they have value as a confidence measure. These use-cases are shown to ultimately be concerned about robustness [chen-mueller-2024-quantifying, macdonald2024generative, zhang2023pacelmpromptingaugmentationcalibrated] which we will highlight. chen-mueller-2024-quantifying develop a Trustworthy Language Model (TLM) where Reliability and explainability is added to every LLM output and decision-making is conducted using trustworthiness scores. This score quantifies the certainty of the model to its answers by combining checking the consistency and the level of verbal numeric confidence of the model’s responses. zhang2023pacelmpromptingaugmentationcalibrated, in the context of cloud incident root cause analysis, develop a method for on-call engineers to make decisions on whether to adopt model predictions using a method that incorporates an LLM-based confidence estimator using verbal confidence due to the black-box nature of the APIs of SOTA models. In macdonald2024generative they build an LLM-based product to structure unstructured data and score customer conversations for developing sales and customer support teams. These scores are based on level of customer satisfaction and given relevant (proprietary) context, the LLM produces a score (verbally) and reasoning for that score. In all of these use-cases it is clear that compromising the robustness of these verbal scores creates a risk of system malfunction or a degradation in operation. Likewise, in pmlr-v235-mohri24a, although not directly related to industry, show the usefulness of utilizing verbal confidence in a specific application, which is that of generating conformal factuality guarantees for language models. They reveal that verbal confidence outperforms the arbitrary confidence baselines (i.e., randomly generated confidence scores and ordinal confidence), demonstrating its value.
@@
Many prior works such as dhuliawala-etal-2023-diachronic, 10.1145/3351095.3372852, 10.1145/3491102.3501967, have established how harmful miscalibrated (skewed) confidence scores are in human studies and downstream human-AI collaboration applications, showing that user behaviour changes greatly in response to AI confidence scores, leading to over reliance when an AI system is overconfident, as one example. Most notably, dhuliawala-etal-2023-diachronic shows in the context of human collaboration with LLMs how un-confidently correct stimuli reduce user trust with only a few miscalibrated instances creating long-term poor performance on tasks.
Moreover, works such as bernsohn-etal-2024-legallens, 50013, 10399826, 10.1136/amiajnl-2011-000291 highlight the impacts of altered confidence scores on downstream authority mechanisms in legal and medical domains, since many such applications that involve tasks such as diagnosis rely on model uncertainty to determine further review. In bernsohn-etal-2024-legallens as an example, in the task of detecting legal violations, entities with high confidence undergo additional processing, and therefore reducing the confidence estimates in the examples would compromise the operation of the system.
@@
In addition to harm done by reduced confidence scores (when answers remain unchanged), our key contribution is also to demonstrate verbal confidence can be used to effectively craft conventional adversarial examples (i.e., ones that lead to answer changes) as well, which presents issues in practical scenarios (e.g., misclassified cases). This is important because such scores are easy to obtain for adversaries for black-box models, and can be used to manipulate the confidence measures that can most easily be asked for and understood by human users.
@@
Our formulation of this issue also has contributions in terms of its implications towards model honesty and how models express their level of knowledge/competence through confidence. A key aspect of this study is to determine how easy their behaviour is to manipulate through various types of mostly superfluous modifications, and through our analysis determining among which aspects there are the most vulnerabilities. In short, one of our objectives is to see how models “perceive” confidence overall and hence we look at multiple different aspects of how models react to different changes of input. We primarily test generalist models to get a broad view of the problem of robustness and to avoid making any conclusions that would only apply to a singular domain.
@@
We study three different threat vectors. An adversary can also target demonstrations or the system prompt too without a user being aware when they send their own unmodified query. This can be a form of data poisoning attack since, for example, adversarially modified demonstrations could be used by a developer of an LLM system unknowingly from a database of seemingly benign demonstrations.
@@ Appendix B Details of Perturbation-based Attacks
The general algorithms for the four different character and word-level adversarial attack methods we cover can be seen in Algorithm 2. For VCA-TF and VCA-TB we make modifications on the black-box algorithms to make the algorithm work in the context of attacking confidence scores generated using CEMs, as these algorithms are originally based on using the predicted softmax class-wise probabilities derived from model logits. Refer to more details in DBLP:conf/aaai/JinJZS20 and DBLP:conf/ndss/LiJDLW19 respectively for more detailed settings as we are faithful to the original algorithm except with respect to the scoring functions used. For the Typos and SubSwapRemove attacks, we make use of our own design choices which we detail below. In the algorithm, the potential options for the target attack text $\mathbf{X}$ are user queries, system prompts, or demonstrations. Additional input text $\mathbf{Q}$ refers to the two remaining text inputs that are not undergoing the attack (e.g., if the user query is targeted, $\mathbf{Q}$ represents the system prompt and demonstration). The parts that are attacked for user queries refer to the question only, without the multiple choice options. For system prompts the entire prompt detailing the task instructions (i.e., how to generate answer and confidence score) can be attacked. For the demonstration attacks, the demonstration is a one-shot example for a given CEM. Within the example, the question, multiple choice options, and the expected answer can be targeted.
For VCA-TB, the possible random bugs that can be generated include inserting spaces into words, deleting characters, swapping adjacent letters, swapping characters with visually similar characters, and substituting words with one of its nearest neighbours. We use WordNet [wordnetcitation] to search for synonyms for each token and calculate the semantic similar using Euclidean distance with GloVe embeddings [pennington-etal-2014-glove]. Likewise, with VCA-TF, we use the same approach to generate a list of synonyms, although we also ensure that synonyms contain the same part of speech (POS) tags using NLTK [https://doi.org/10.48550/arxiv.cs/0205028].
To perform the Typos attack, we generate a series of common typos in a sentence (GenerateCommonTypos function). These typos include substitutions with nearby characters on an English keyboard, random character deletions, and adjacent character swaps. A probability of $0.1$ is used to select the percentage of characters in a full text sequence to modify each iteration. We cap the number of iterations at 25.
@@
1:Input: Confidence Estimation Method $z$ for LLM, Tokenized Target Text $\mathbf{X}$, Additional Input Text $\mathbf{Q}$ Sentence Similarity Function SIM(), Perturbation Similarity Rejection Threshold $\epsilon$, Max Iterations $\tau$, # Synonym Candidates $N_{syn}$
@@
[…]
@@
[…]
@@
15:  BugBug $\leftarrow$ Randomly Selected TextBugger Bug on $w_{i}$
16:  $\mathbf{\hat{X}}_{temp}$ $\leftarrow$ $\mathbf{\hat{X}}_{w_{i}<>Bug}$ $\triangleright$ Bug replaces original token in $\mathbf{\hat{X}}$
17:  if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
@@
19:     $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
@@
27:  for all $\mbox{cnd}\in\mbox{Candidates}$ do
28:   $\mathbf{\hat{X}}_{temp}$ $\leftarrow$ $\mathbf{\hat{X}}_{w_{i}<>cnd}$ $\triangleright$ cnd replaces $w_{i}$ in $\mathbf{\hat{X}}$
29:   if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
@@
31:      $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
32:      topscore == $z(\mathbf{Q},\mathbf{\hat{X}}_{temp})$
@@
43:     $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
@@
51:  MOD $\leftarrow$ Randomly Choose Perturbation Type from {Sub, Swap, Remove}
@@
53:  if SIM($\mathbf{X},\mathbf{\hat{X}}_{temp})>\epsilon$ then
@@
55:     $\mathbf{\hat{X}} \leftarrow$ $\mathbf{\hat{X}}_{temp}$
@@ Appendix C Detailed Experimental Settings
Dataset Details: (1) MedMCQA [pmlr-v174-pal22a] is a multiple-choice question answering (MCQA) dataset consisting of 2816 questions with four choices that are based on medical entrance exam questions. (2) TruthfulQA [lin2022truthfulqa] is a MCQA dataset consisting of questions that cover a broad range of safety-critical topics like health, law, finance and politics and is designed so that humans can easily answer incorrectly due to misconceptions. We use the MC1 (Single-true) subtask, and only use multiple choice questions with 4 total answers for consistency yielding 203 total examples. (3) StrategyQA [geva2021strategyqa] consists of 2,780 true or false open-domain, general topic questions that require implicit reasoning steps to answer. Lastly, we use SST-2 [socher-etal-2013-recursive] from the popular AdvGLUE [wang2021adversarial] adversarial benchmark for our baseline confidence attack agnostic results. In the benchmark, SST-2 consists of 137 movie reviews annotated for binary (positive/negative) sentiment. Instances in AdvGLUE underwent a comprehensive series of word-level, sentence-level, adversarial algorithmic modifications, in addition to human-generated perturbations to most effectively deceive a wide range of language models. For our experiments, we use the development sets for MedMCQA and TruthfulQA, and the training set for StrategyQA.
Model Details: Llama-3-8B [llama3modelcard] and GPT3.5-turbo [OpenAI2024] are primarily used for evaluation, with GPT-4o [OpenAI2024] and Llama-3-70B [llama3modelcard] being utilized for certain additional results. The Llama models are ran locally while the GPT models are ran through the OpenAI API. In terms of sampling parameters for each model, we utilize a temperature of 0.0 for Base, CoT, and MS method, along with a top_p of 0.95. For SC and SC-CA, we set the temperature to 0.7 and sample three different responses and average the confidence scores. We limit the number of generated tokens to 400. For the MS method, we calculate the geometric mean of the generated confidence scores for each step including the final.
Sentence Similarity: To ensure adversarial examples preserve the original meaning, we encode both the original and adversarial example using Universal Sentence Encoder (USE) [cer2018universal]. We then calculate cosine similarity between the vectors and forgo perturbations that cause a fall below a threshold of 80% similarity to keep adversarial examples retaining similar meaning of the original sentences similar to the approach in wang2023adversarial.
Confidence Extraction: Answers and confidence are extracted from the generated answer of the LLM using robust regular expressions looking for all possible mentions of letter answers and their corresponding numeric confidence. In the rare cases where models refuse to answer or provide a confidence score (or the regular expression fails), the final answer defaults to none-of-the-above and the final confidence score is set to be one with max entropy possible (i.e., $1/(\mbox{Number of Multiple Choice Options})$) respectively. Over the course of optimizing an attack, if no confidence is provided, the perturbation is rejected unless the answer choice flips as well (signifying a successful attack).
Table 7: Description of SC-CA CEM.
CEMs Description Example Answer
SC-CA Think step by step, produce letter answer and confidence at the end. Sample and average multiple answers. (1) … therefore the answer is B, 90% (2) I believe … so the answer is B, 95% (3) …, the answer must be A, 80%. Final Answer B, 88.33%